OpenStudy (anonymous):

solving system of equations by substitution. y=6x - 11 -2x - 3y = -7

OpenStudy (anonymous):

-2x - 3(6x-11) = -7 -2x -18x + 33 = -7 -20x = -7 - 33 -20x = -40 x = 2

OpenStudy (anonymous):

you just put y from the first equation into the second equation

OpenStudy (anonymous):

How is that? @wowgirl111

OpenStudy (anonymous):

its good! thanks(: